OpenStudy (anonymous):

i need work shown write .569 as a percent

OpenStudy (rishabh.mission):

http://www.wolframalpha.com/input/?i=i+need+work+shown+write+.569+as+a+percent mb its can help u

OpenStudy (mathstudent55):

A percent is a ratio of a number to 100 0.569 is read as five-hundred sixty-nine thousandths, so as a fraction it is written as 569/1000. All you are trying to find out is 569/1000 is how many hundredths: 569/1000 = x/100 Cross multiply: 1000x = 100*569 1000x = 56,900 x = 56.9 Answer: 0.569 = 56.9%